A valószínűségszámítás I. vizsgaeredményei a 2002–2003 tanév második félévében.

diákok:  jegy  időpont
Aradi Zsolt  4, jó  június 27.
Ács Gábor  2, elégséges  május 30.
Bálint Gergely  3, közepes  június 27.
Balogh Erika  2, elégséges  május 30.
Barabás Timea  2, elégséges  augusztus 27.
Baráth Terézia  2, elégséges  június 12.
Barna Marianna  2, elégséges  augusztus 27.
Barczi Krisztina  3, közepes  június 27.
Bazsó András  5, jeles  június 12.
Bulátki János  3, közepes  június 27.
Buzogány Ágota  5, jeles  június 27.
Ceglédi Mariann  3, közepes  június 27.
Csüllög Irén  4, jó  május 26.
Dascál Judit  3, közepes  június 27.
Elek Judit  5, jeles  május 30.
Erdős Katalin  3, közepes  június 27.
Fazekas Gabriella  5, jeles  június 27.
Galiné, Mátyás Veronika  3, közepes  június 27.
Gyöngy Dániel  3, közepes  június 27.
Hellebrand Márk  2, elégséges  június 27.
Hepp Róbert  2, elégséges  május 30.
Hubert Balázs  3, közepes  június 12.
Izsó Mónika  4, jó  június 27.
Jáger Erzsébet  4, jó  június 27.
Kalapos Adrienn  4, jó  június 27.
Karakas Éva  3, közepes  június 27.
Károlyi Zsuzsanna  3, közepes  június 27.
Katona Kornél  2, elégséges  június 12.
Katonka Pál  2, elégséges  június 12.
Kiss István  2, elégséges  június 27.
Kolbe Katalin  2, elégséges  augusztus 27.
Kovács Szilvia  2, elégséges  augusztus 27.
Lévai Margit Tünde  3, közepes  június 27.
Mártha Lívia  2, elégséges  június 27.
Mező István  4, jó  június 12.
Molnár Kinga  3, közepes  május 30.
Nagy Adrienn  2, elégséges  augusztus 27.
Nagy Csilla  2, elégséges  június 27.
Nemes Melinda  5, jeles  június 27.
Pályi Zsanett  3, közepes  június 27.

Pászner Marietta  5, jeles  június 27.
Pék Johanna  4, jó  május 30.
Piricsi Renáta  4, jó  június 27.
Spekker Melinda  3, közepes  június 6.
Süli–Zakar Ákos  3, közepes  június 27.
Székely Eszter  2, elégséges  június 27.
Szilágyi Renáta  3, közepes  június 27.
Szűcs Orsolya  4, jó  május 30.
Takács Margit  3, közepes  május 30.
Teitelbaum Zsuzsanna  3, közepes  május 30.
Tímár Olga  3, közepes  május 30.
Tolnai Eszter  3, közepes  május 30.
Tóth György  2, elégséges  augusztus 27.
Tóth Krisztina  2, elégséges  június 27.
Varga Valéria  2, elégséges  június 27.
Veres Judit  4, jó  május 30.
Veres Rosina  4, jó  május 30.
Zachár Timea  2, elégséges  június 12.
Zám Katalin  2, elégséges  június 27.

Major Péter